OpenStudy (anonymous):

Which fraction is halfway between and 3/5and 7/10? 13/10 13/20 7/50 7/15

OpenStudy (imnotgoodatmath):

i think 13/20

OpenStudy (anonymous):

3/5=.6 7/10=.7 We need to find the fraction that is equal to .65 13/10=1.3 7/50=.14 7/15=about .5 Therefore, the answer is 13.20, which is .65.